OpenStudy (anonymous):

How do i solve squareroot 2 sinx+1=0

OpenStudy (anonymous):

is the 2sinx+1 all under the radical and then it's = to 0 ?

OpenStudy (anonymous):

oh sorry just the 2 is under the radical

OpenStudy (anonymous):

ok you can isolate the radical by moving everything to the other side. then raise both sides to the power of 2. But i am not sure what varible you are solving for

OpenStudy (anonymous):

but if sin is the coeffeciant of x then you should be able to divide to get them brake them down but i am not sure of this

OpenStudy (anonymous):

oh ok thank you anyway

OpenStudy (anonymous):

\[\sqrt{2}\sin x =-1\]\[\sin x =-1/\sqrt{2}\]\[x =\sin^{-1} -1/\sqrt{2}\]\[x =\pi/4\]

OpenStudy (anonymous):

I'm sorry\[x =5\pi/4\]

OpenStudy (l):

sqrt_2 . sin_x + 1 = 0 sin_x = -1/sqrt_2 x = -pi/4 or 5pi/4 *